# Tug-of-war between idiom’s figurative and literal meanings in LLMs

### v3

@@ Appendix C Layer-wise interpretation flow through activation patching
Figure 9: Description of the activation patching experiment for identifying interpretation flow. L represents literal interpretation; F represents figurative interpretation. (a) By replacing ‘because’ token $s_{f}$ activation with $s_{a}$, we expect literal interpretation to be added in the intermediate channel while keeping the direct route flow to be activated with naive figurative meaning. (b) By replacing ‘because’ token $s_{l}$ activation with $s_{a}$, we expect figurative interpretation to be added via intermediate path, while keeping both paths for literal interpretation.
@@
Figure 10: Layer-wise interpretation shift $\Delta I(s_{\mathrm{x}}\hookleftarrow s_{a}),\,\mathrm{x}\in\{f,l\}$ after activation patching on the ‘because’ token across three patch directions. Y-axis: Mean values of $\Delta L(s_{\mathrm{x}}\hookleftarrow s_{a})$, $\Delta F(s_{\mathrm{x}}\hookleftarrow s_{a})$. X-axis: Layers. Gray dashed line: $\Delta I=0$ (no effect). Red asterisk (*): Significant shift between literal ($C_{l}$) and figurative ($C_{f}$) candidate probability (paired $t$-test, $p<0.05$).
As expected, by adding idiom’s interpretation via intermediate channel, since it’s adding literal interpretation, there’s a increase in $\Delta L(x_{f}\hookleftarrow s_{a})$ by $0.1$, whereas $\Delta F(x_{f}\hookleftarrow s_{a})$ drops by $-0.07$ across layers (Figure 10(a)). Whereas, when adding the idiom’s figurative interpretation into intermediate path, there’s a big increase in $\Delta F(s_{l}\hookleftarrow s_{a})$ by $0.26$ at layer 7 and decrease in $\Delta L(s_{l}\hookleftarrow s_{a})$ by $-0.17$. This implies that when there’s idiomatic representation passing the intermediate path, to make final prediction as close to figurative one, the model tries to amplify figurative meaning that is passing the intermediate path so that the total amount of activations for figurative interpretation to be same, which should be more than its literal one.
That is, injecting the idiom’s literal activation into the intermediate channel increases $\Delta L(x_{f}\hookleftarrow s_{a})$ by up to $0.1$ in the mid-layers (4-9). Conversely, injecting figurative activation raises $\Delta F(s_{l}\hookleftarrow s_{a})$ by $0.26$. This asymmetric increase indicates that when an idiomatic representation passes the intermediate path, the model actively amplifies figurative interpretation so it dominates the final prediction.
Appendix D Analyzing edge-level intermediate channel through attention knockout
To trace the flow of information, we perform fine-grained intervention on the MHSA sublayers, since they mediate communication between token positions, therefore, any critical information must pass through them Geva et al. (2023).
To locate the intermediate and direct path for each interpretation, we focus on the attention scores between idiom to subsequent token (i.e., because) for the intermediate path, and the idiom to last token of the sentence for the direct path findings. We investigate how this interpretation of flow is done internally by knockout the attention edges between target tokens. When critical attention edges for a certain interpretation of the idiom are blocked, it will result in severe degradation in probabilities of certain candidates. Therefore, we test whether critical information propagates between two hidden representations at a specific layer, by zeroing-out all the attention edges between them.
Formally, let $i,j\in\{1,\dots,T\}$ with $j\leq i$ be two positions in the sequence. To block token $i$ (query) from attending to token $j$ (key) at layer $\ell<L$, we overwrite the corresponding entries in the causal self-attention mask $M$ at layer $\ell+1$ across all heads $h\in\{1,\dots,H\}$:
$M_{i,\,j}^{\ell+1,\,h}\;=\;-\,\infty\quad\forall\,h\in\{1,\dots,H\}.$
(a) Idiom $\rightarrow$ Subsequent
(b) Idiom $\rightarrow$ Last
@@ Appendix D Analyzing edge-level intermediate channel through attention knockout
Figure 11: Layer-wise interpretation shift $\Delta I(s_{a})$ after knock out the attention edges between tokens. Y-axis: Mean values of $\Delta L(s_{a})$, $\Delta F(s_{a})$. X-axis: Layers. Gray dashed line: $\Delta I=0$ (no effect). Red asterisk (*): Significant shift between literal ($C_{l}$) and figurative ($C_{f}$) candidate cumulative probabilities (paired $t$-test, $p<0.05$).
This restricts the query position from obtaining information from the key position, at that particular layer ($\ell+1$). For each layer, we block attention edges between (a) idiom span tokens $\rightarrow$ subsequent token (b) idiom span tokens $\rightarrow$ last token. If the specific path is important for a specific interpretation, then there should be significant drop of $I(s_{a})$ value. Figure 11 shows the result.
In Figure 11(a) at layer 5, it shows that there’s a drop in the $\Delta F(s_{a})$ by $-0.02$, while increase in $\Delta L(s_{a})$, by $0.02$. This shows that ablating attention from the idiom span to ‘because’ token disrupts the indirect path that would do disambiguation process where it suppresses the literal one. Without the idiom attending to the ‘because’ token, the model can no longer effectively disambiguate figurative meaning from literal meaning.
At layer 11 and 13, as in Figure 11(b), $\Delta L(s_{a})$ declines while $\Delta F(s_{a})$ rises. This divergence indicates that ablating the direct idiom to last token pathway prevents the model from passing the literal interpretation.
@@
She was looking for a needle in a haystack because she was a She was searching for a neddle within a straw because she was a She was facing an impossible search because she was a needle, craft, seam, straw, tiny, hay, detective, master, haystack, witch, farmer, pro, camel, professional, busy, crazy, giant, chicken, cow, neat woman, victim, single, young, black, female, girl, slave, mother, stranger, new, first, non, private, white, novice, minority, prisoner, foreign, mom

### v4

@@ Appendix E Activation patching for information flow with other models
Llama3.1-8B
[…]
Qwen2.5-7B
[…]
Figure 12: Layer-wise interpretation shift after patching in activations from $s_{f}$ and $s_{l}$ and vice versa. The red asterisk (*) marks layers where the difference between $\Delta L$ and $\Delta F$ exceeds the average difference across layers (paired $t$-test, $p<0.05$).
Appendix F Knockout for context disambiguation with other models
Across different models, Figure 13 consistently indicate that early MLP layers retrieve idioms’ figurative meanings, and following layers promotes contextual cues via MHSA when a literal context is present.
Qwen2.5-0.5B
(a) MLP FC
(b) MHSA FC
(c) MLP LC
(d) MHSA LC
Llama3.1-8B
(e) MLP FC
(f) MHSA FC
(g) MLP LC
(h) MHSA LC
Qwen2.5-7B
(i) MLP FC
(j) MHSA FC
(k) MLP LC
(l) MHSA LC
@@ Appendix F Knockout for context disambiguation with other models
Figure 13: Sublayer-wise interpretation shift $\Delta I(C+s_{a})$ after ablating activations at idiom span, for contexts $C\in\{FC,LC\}$. Y-axis: Mean values of $\Delta L(C+s_{a})$, $\Delta F(C+s_{a})$ with 95% confidence intervals. X-axis: Layers. Gray dashed line: $\Delta I=0$ (no effect). Red asterisk (*): Significant difference between $\Delta F(C+s_{a})$ and the others (paired $t$-test, $p<0.05$). The difference at * marked layer is larger than the average difference across all layers.
@@
She was looking for a needle in a haystack because she was a She was searching for a needle within a straw because she was a She was facing an impossible search because she was a needle, craft, seam, straw, tiny, hay, detective, master, haystack, witch, farmer, pro, camel, professional, busy, crazy, giant, chicken, cow, neat woman, victim, single, young, black, female, girl, slave, mother, stranger, new, first, non, private, white, novice, minority, prisoner, foreign, mom